oHE of these 7. In a survey of 600 people with 317 claiming they would vote yes, find and interpret the 9/% C.I that the proportion of yes voters is 5283 b) we are sure of 317 yes votes c) we know that the proportion of yes voters lies within (4758, .5808) d) we are 99% confident that the proportion of yes voters is .5283 e) we are 99% confident that the proportion of yes voters lies within (4758,5808) 8. The area of each critical region for a two tailed test at a .01 is: a) .005 b) .01 c) .025 d) .05 e) none of these The critical z-scores for one tai a) +1.645 b) 1.96 c) 2.33 d) +2.58 e) none of these 9. led tests at a01 are: The critical t-scores for one tailed tests at α .05 with 12 for sample size are: a) t2.18 b) t1.78 c) 2.20 d) t1.80 e) none of these 10. 11. The form(s) for the alternative hypothesis is/are a) H : μ=a#, b) H : μ<a#” c) H: μ>a#, d) Ha: μs# e) all ofthese except (a) luiut thats wrong so you
Answer #1

Ans:

7)

sample proportion=317/600=0.5283

99% confidence interval for p

=0.5283+/-2.58*sqrt(0.5283*(1-0.5283)/600)

=0.5283+/-0.0526

=(0.4758, 0.5809)

option e is correct.

8)

alpha=0.01

alpha/2=0.01/2=0.005

9)

alpha=0.01

critical z value(one tailed)=-2.33 or 2.33

option c is correct.

10)n=12

df=12-1=11

alpha=0.05

critical t value(one tailed)=-1.80 or 1.80

option d is correct.

11)option e is correct(except a)